# Physics-informed neural networks for modeling dynamic linear elasticity

### v2

@@ 3 Linear elasticity
where the scaling factors $\bm{\cdot}_{c}$ are chosen to be the absolute maximum value of the quantities under consideration for the forward problem. We assume $x_{c}=y_{c}=l_{c}$ and $\lambda_{c}=\mu_{c}$. For forward simulations, we know the value of the material properties. However, for the inverse problem or surrogate modeling, since we do not have any prior information on the material properties, we have to assume a value for $\lambda_{c}$ and $\mu_{c}$. In a practical scenario, for an inverse problem, for training data, one will perform measurements on a newly manufactured material and one would know a rough value for the properties which can then be chosen for the scaling. Also, while performing neural network prediction, the quantities with $\bm{\cdot}^{*}$ will be scaled back to original range.
@@ 4 PINN model for linear elasticity
The norm .$\|.\|_{data}=\frac{1}{N_{data}}\sum_{n=1}^{N_{data}}|{q}({\bf x}^{d}_{n},t^{d}_{n})|^{2}$ where $q$ is a quantity of interest. Likewise, .$\|.\|_{eqn}=\frac{1}{N_{eqn}}\sum_{n=1}^{N_{eqn}}|{q}({\bf x}^{e}_{n},t^{e}_{n})|^{2}$. Also, $\alpha_{1}$ to $\alpha_{5}$ in front of the terms in Eq. 34 are the weight coefficients. For the neural network training, we consider couple of useful suggestions provided by (Wang et al., 2023). First, we use the modified neural network architecture proposed by (Wang et al., 2021) where the forward pass of a $L$-layer architecture is defined as
@@ 5.1 PINN model for forward problem
Figure 3: Total loss for different number of neurons and layers with $6.76$ training data.
Figure 4: Total loss for different number of neurons and layers with $11.52$ training data.
Figure 5: Total loss for different number of neurons and layers with $23.9$ training data.
@@ 5.1.1 Hyperparameter search
For all our hyperparameter sweeps, we train the model for a total of $6000$ epochs. For the first $2000$ epochs, we use a learning rate of $\eta=0.001$. For the next $2000$ epochs, we keep $\eta=0.0001$ and for the last $2000$ epochs, $\eta=0.00001$. As for the activation functions we use a $\tanh$ function for all the layers except for the last layer where we use a linear activation function (Raissi et al., 2019). The training is performed using a NVIDIA Tesla V100 graphic card. We choose three sets of neurons $(64,128,256)$ per hidden layer and number of hidden layers $(4,5,6)$ for three training data sets. For the same, the total loss curves are shown in Figs. 3, 4 and 5. From Fig. 3, for the training data of $6.76$ with batch size of 142, we find the computationally optimal set of neurons and layers to be $(256,4)$. Similarly, from Fig. 4, for $11.52$ training data with batch size of 121, the best set of neurons and layers are $(128,4)$. Finally, from Fig. 5, for $23.9$ training data with batch size of 251, it is $(128,5)$. Here the % data is shown with respect to boundary data from the FEM solution. Also, we use $500$ random collocation points for the calculation of $\mathcal{L}_{eqn}$ at each gradient descent step. These collocation points are sampled using Latin hypercube sampling (LHS) method (McKay et al., 1979).
@@
$256$ $4$ $6.76$ $500$ $3.12\times 10^{-3}$ $5.12\times 10^{-4}$
$128$ $4$ $11.56$ $500$ $3.07\times 10^{-3}$ $5.48\times 10^{-4}$
$128$ $5$ $23.9$ $500$ $2.91\times 10^{-3}$ $5.2\times 10^{-4}$
@@ 5.1.2 Comparison with FEM solution
For the best of hyperparameter sets for their corresponding training data %, Table 1 shows the NRMSE error $\epsilon$ in $u_{x}$ and $u_{y}$ as compared with the FEM solution. We compute $\epsilon$ at 101 $\times$ 11 $\times$ 100 spatio-temporal points which corresponds to 101 $\times$ 11 sparse spatial grid points and 100 time instants. We observe that for all three training data sets, the testing space error is of the order $10^{-3}$ and $10^{-4}$ for the displacements $u_{x}$ and $u_{y}$ which is reasonably small. Errors for training data of $6.76$ is of the same order as that of other training data sets. So, for the following results in this section we have chosen the training data of $6.76$.
@@
Next, we show the spatial NRMSE error $\epsilon$ in displacements and stresses with respect to time in Fig. 8. We observe that error is low and there is no general trend of error growth or decay in time. Considering we have shown convincing results from PINN for a sparse training data set of $6.76$, if we analogously assume that we have a sparse set of measurement data, we could use this data along with a PINN model to obtain complete information of the displacements and stresses for all assumed spatio-temporal points for this dynamic system. This makes practical sense because we usually perform measurements at limited sensor locations due to cost considerations. Another advantage of the PINN model is that we get solution information at the interior of the structure which has similar accuracy to the FEM solution.
@@
Figure 7: Line plots of the quantities of interest at $(x=120\ \text{mm},\ y=8\ \text{mm},\ t=1.72\ \text{s})$ for material properties $\lambda=0.533334$ MPa and $\mu=0.1$ MPa. Plots in (a) to (e) show the comparison of displacements and stresses obtained by PINN with respect to the FEM solutions.
@@ 5.2 PINN model for inverse problem
In this section, we discuss the results obtained using PINN in the context of inverse modeling where one knows the experimental data and wants to find the material properties ($\lambda,\mu$). Here we use FEM simulation data as an analogue of experimental data. The problem definition is same as shown in Fig. 2 except here we apply an oscillating excitation of $-\sin(2\text{t})$. Also, in this case, we consider the material as steel which has a density of $7.85\times 10^{-6}$ $kg/mm^{3}$ and it has the material properties $\lambda=115385$ MPa and $\mu=76923$ MPa. We run FEM simulation for $3.11$ seconds. For training the PINN model, we use all data from the boundaries as well as sparse data from interior of the domain which accounts for $4.4$ of entire simulation data. Also, temporally, we use the data at all the intermediate times which are available. As for the hyperparameters, we use $256$ neurons per hidden layer with $4$ hidden layers, and $1000$ random collocation points at each gradient descent step. Since we are interested in learning the material parameters, we consider these as trainable parameters along with weights and biases of the neural network. We consider $N_{\lambda}$, $N_{\mu}$ as trainable parameters where $\lambda^{*}=f(N_{\lambda}),\mu^{*}=f(N_{\mu})$. The choice of function $f$ significantly affects the learning process. Once the model is trained we can rescale these parameters to get actual material parameters by $\lambda=\lambda_{c}\lambda^{*}$ and $\mu=\mu_{c}\mu^{*}$.
@@
We train the PINN model for first 1000 epochs with learning rate $\eta=0.001$ and next 10000 epochs with $\eta=0.0001$. The total training data is 4.4 % of entire simulation data. We have considered batch size of 225 for the training. We show in Fig. 9 the total loss and the evolution of the predicted parameters $\lambda$ and $\mu$ over the course of training epochs. We observe that as the training progresses, the parameters tend to reach closer to their true value which is shown as a horizontal line. We get the PINN prediction error for $\lambda$ and $\mu$ to be $2.57$ and $1.9$ respectively. Although we find the prospect of PINN model’s use in inverse modeling exciting, in this particular application, we find its practical use to be limited in the industrial context. One reason for this is that in most industrial settings, stress measurements inside the material which is required for training is not straightforward, especially if the geometry is complicated. Secondly, it can result in increased experimental costs because as more data is needed, more sensors are required. Due to these reasons, data quantity and quality for training the PINN model in this context will be limited. Hence, we will not expand upon the results of inverse models further in this study. However, the results we obtained in this context could be exciting for applications based on tomography.
@@ 5.3 PINN based surrogate model for material property estimation
Due to practical constraints on data procurement via experimental measurements, traditionally, numerical simulations are a mainstay for problems in solid mechanics. However, in this context of material property estimation, one resorts to performing repeated simulations for various iterative combinations of the parameters. This iterative search ends when one finds the parameters for which the experimental data matches with the simulation. Thus, one incurs high amount of computational cost. In this section, we will show how a PINN model can be used to circumvent this issue. We consider the same problem of a beam with an oscillating excitation at the free end as shown in Fig. 2. However, in this case, we train the PINN model with simulation data of six different values of the material property $\mu$ while keeping $\lambda=0.533334$ MPa constant. The parameter $\mu$ is also given as additional input feature to the network along with spatial and temporal coordinates as $(x,y,t,\mu)$. The training values chosen for $\mu$ are shown in the six five rows of Table 3. In the same table, the bottom three rows below a horizontal line denote the testing space $\mu$ values.For training, in space, we choose $23.9$ of the boundaries data and in time we choose all time instances of FEM simulation. This training data accounts for nearly $0.98$ of entire simulation data. We try to constrain model toward learning solution based on equations properly, in order to fulfill that we give more weightage to the equation loss by selecting $\alpha_{1}=\alpha_{2}=10^{3}$ so as to make equation loss tends towards even lower order of magnitudes. Also, we show in Table 3 the error values $\epsilon$ for $u_{x}$ and $u_{y}$. We observe that the testing errors are one order higher than the training errors. However, they are still low enough to obtain physically meaningful solutions. In Fig. 10, we show the contour plots from the PINN model and FEM for testing $\mu=0.075$ MPa. We notice a high accuracy in the prediction of $u_{x}$ and $u_{y}$ from PINN. Even if stresses are not exactly matching to the FEM solution, they are of the same order of magnitude. Furthermore, in Fig. 11, we show the error $\epsilon$ as a function of time for a particular location in space. We observe again a very close match between PINN and FEM for the displacements. Barring the training time for PINN, the prediction based on PINN for new parameters $\mu$ is extremely fast, usually seconds. For the same problem where $\mu=0.075$ MPa, the FEM simulation takes approximately $20$ minutes on a standalone Intel Core i7 CPU. Hence, the PINN model easily outperforms the FEM simulation in terms of computational cost while maintaining reasonable accuracy.
@@
Figure 11: Line plots of the quantities of interest at $(x=120\ \text{mm},\ y=8\ \text{mm},\ t=1.72\ \text{s})$ for material property $\mu=0.075$ MPa. Plots in (a) to (e) show the comparison of displacements and stresses obtained by PINN with respect to the FEM solutions.
@@ 5.4.2 Results and discussion (3D)
Figure 13: Contour plots of the quantities of interest in the $xy$-plane at $z=5\ mm$ and $t=2.29$ s for the material property $\mu=0.075$ MPa. Plots in (a) denote FEM solutions. Plots in (b) denote PINN solutions. Plots in (c) denotes the absolute error.
Figure 14: Line plots of the quantities of interest at $(x=120\ \text{mm},\ y=8\ \text{mm},\ z=5\ \text{mm},\ t=2.29\ \text{s})$ for material property $\mu=0.075$ MPa. Plots in (a) to (e) show the comparison of displacements and stresses obtained by PINN with respect to the FEM solutions.
@@
$0.2$ $3.5\times 10^{-3}$ $1\times 10^{-3}$ $1.1\times 10^{-3}$

### v3

@@ 3 Linear elasticity
[…]
@@ 4 PINN model for linear elasticity
The norm $\|.\|_{data}=\frac{1}{N_{data}}\sum_{n=1}^{N_{data}}|{q}({\bf x}^{d}_{n},t^{d}_{n})|^{2}$ where $q$ is a quantity of interest. Likewise, $\|.\|_{eqn}=\frac{1}{N_{eqn}}\sum_{n=1}^{N_{eqn}}|{q}({\bf x}^{e}_{n},t^{e}_{n})|^{2}$. Also, $\alpha_{1}$ to $\alpha_{5}$ in front of the terms in Eq. 34 are the weight coefficients. For the neural network training, we consider couple of useful suggestions provided by (Wang et al., 2023). First, we use the modified neural network architecture proposed by (Wang et al., 2021) where the forward pass of a $L$-layer architecture is defined as
@@ 5.1 PINN model for forward problem
Figure 3: Total loss for different number of neurons and layers with $6.76$ % training data.
Figure 4: Total loss for different number of neurons and layers with $11.52$ % training data.
Figure 5: Total loss for different number of neurons and layers with $23.9$ % training data.
@@ 5.1.1 Hyperparameter search
For all our hyperparameter sweeps, we train the model for a total of $6000$ epochs. For the first $2000$ epochs, we use a learning rate of $\eta=0.001$. For the next $2000$ epochs, we keep $\eta=0.0001$ and for the last $2000$ epochs, $\eta=0.00001$. As for the activation functions we use a $\tanh$ function for all the layers except for the last layer where we use a linear activation function (Raissi et al., 2019). The training is performed using a NVIDIA Tesla V100 graphic card. We choose three sets of neurons $(64,128,256)$ per hidden layer and number of hidden layers $(4,5,6)$ for three training data sets. For the same, the total loss curves are shown in Figs. 3, 4 and 5. From Fig. 3, for the training data of $6.76$ % with batch size of 142, we find the computationally optimal set of neurons and layers to be $(256,4)$. Similarly, from Fig. 4, for $11.52$ % training data with batch size of 121, the best set of neurons and layers are $(128,4)$. Finally, from Fig. 5, for $23.9$ % training data with batch size of 251, it is $(128,5)$. Here the % data is shown with respect to boundary data from the FEM solution. Also, we use $500$ random collocation points for the calculation of $\mathcal{L}_{eqn}$ at each gradient descent step. These collocation points are sampled using Latin hypercube sampling (LHS) method (McKay et al., 1979).
@@
$256$ 44 $6.76$ % $500$ $3.12\times 10^{-3}$ $5.12\times 10^{-4}$
$128$ 44 $11.56$ % $500$ $3.07\times 10^{-3}$ $5.48\times 10^{-4}$
$128$ 55 $23.9$ % $500$ $2.91\times 10^{-3}$ $5.2\times 10^{-4}$
@@ 5.1.2 Comparison with FEM solution
For the best of hyperparameter sets for their corresponding training data %, Table 1 shows the NRMSE error $\epsilon$ in $u_{x}$ and $u_{y}$ as compared with the FEM solution. We compute $\epsilon$ at 101 $\times$ 11 $\times$ 100 spatio-temporal points which corresponds to 101 $\times$ 11 sparse spatial grid points and 100 time instants. We observe that for all three training data sets, the testing space error is of the order $10^{-3}$ and $10^{-4}$ for the displacements $u_{x}$ and $u_{y}$ which is reasonably small. Errors for training data of $6.76$ % is of the same order as that of other training data sets. So, for the following results in this section we have chosen the training data of $6.76$ %.
@@
Next, we show the spatial NRMSE error $\epsilon$ in displacements and stresses with respect to time in Fig. 8. We observe that error is low and there is no general trend of error growth or decay in time. Considering we have shown convincing results from PINN for a sparse training data set of $6.76$ %, if we analogously assume that we have a sparse set of measurement data, we could use this data along with a PINN model to obtain complete information of the displacements and stresses for all assumed spatio-temporal points for this dynamic system. This makes practical sense because we usually perform measurements at limited sensor locations due to cost considerations. Another advantage of the PINN model is that we get solution information at the interior of the structure which has similar accuracy to the FEM solution.
@@
Figure 7: Line plots of the quantities of interest at (x=120mm,y=8mm)(x=120\ $\text${mm},\ y=8\ $\text${mm}) for material properties $\lambda=0.533334$ MPa and $\mu=0.1$ MPa. Plots in (a) to (e) show the comparison of displacements and stresses obtained by PINN with respect to the FEM solutions.
@@ 5.2 PINN model for inverse problem
In this section, we discuss the results obtained using PINN in the context of inverse modeling where one knows the experimental data and wants to find the material properties ($\lambda,\mu$). Here we use FEM simulation data as an analogue of experimental data. The problem definition is same as shown in Fig. 2 except here we apply an oscillating excitation of $\sin(2\text{t})$. Also, in this case, we consider the material as steel which has a density of $7.85\times 10^{-6}$ $kg/mm^{3}$ and it has the material properties $\lambda=115385$ MPa and $\mu=76923$ MPa. We run FEM simulation for $3.11$ seconds. For training the PINN model, we use all data from the boundaries as well as sparse data from interior of the domain which accounts for $4.4$ % of entire simulation data. Also, temporally, we use the data at all the intermediate times which are available. As for the hyperparameters, we use $256$ neurons per hidden layer with $4$ hidden layers, and $1000$ random collocation points at each gradient descent step. Since we are interested in learning the material parameters, we consider these as trainable parameters along with weights and biases of the neural network. We consider $N_{\lambda}$, $N_{\mu}$ as trainable parameters where $\lambda^{*}=f(N_{\lambda}),\mu^{*}=f(N_{\mu})$. The choice of function $f$ significantly affects the learning process. Once the model is trained we can rescale these parameters to get actual material parameters by $\lambda=\lambda_{c}\lambda^{*}$ and $\mu=\mu_{c}\mu^{*}$.
@@
We train the PINN model for first 1000 epochs with learning rate $\eta=0.001$ and next 10000 epochs with $\eta=0.0001$. The total training data is 4.4 % of entire simulation data. We have considered batch size of 225 for the training. We show in Fig. 9 the total loss and the evolution of the predicted parameters $\lambda$ and $\mu$ over the course of training epochs. We observe that as the training progresses, the parameters tend to reach closer to their true value which is shown as a horizontal line. We get the PINN prediction error for $\lambda$ and $\mu$ to be $2.57$ % and $1.9$ % respectively. Although we find the prospect of PINN model’s use in inverse modeling exciting, in this particular application, we find its practical use to be limited in the industrial context. One reason for this is that in most industrial settings, stress measurements inside the material which is required for training is not straightforward, especially if the geometry is complicated. Secondly, it can result in increased experimental costs because as more data is needed, more sensors are required. Due to these reasons, data quantity and quality for training the PINN model in this context will be limited. Hence, we will not expand upon the results of inverse models further in this study. However, the results we obtained in this context could be exciting for applications based on tomography.
@@ 5.3 PINN based surrogate model for material property estimation
Due to practical constraints on data procurement via experimental measurements, traditionally, numerical simulations are a mainstay for problems in solid mechanics. However, in this context of material property estimation, one resorts to performing repeated simulations for various iterative combinations of the parameters. This iterative search ends when one finds the parameters for which the experimental data matches with the simulation. Thus, one incurs high amount of computational cost. In this section, we will show how a PINN model can be used to circumvent this issue. We consider the same problem of a beam with an oscillating excitation at the free end as shown in Fig. 2. However, in this case, we train the PINN model with simulation data of six different values of the material property $\mu$ while keeping $\lambda=0.533334$ MPa constant. The parameter $\mu$ is also given as additional input feature to the network along with spatial and temporal coordinates as $(x,y,t,\mu)$. The training values chosen for $\mu$ are shown in the six five rows of Table 3. In the same table, the bottom three rows below a horizontal line denote the testing space $\mu$ values.For training, in space, we choose $23.9$ % of the boundaries data and in time we choose all time instances of FEM simulation. This training data accounts for nearly $0.98$ % of entire simulation data. We try to constrain model toward learning solution based on equations properly, in order to fulfill that we give more weightage to the equation loss by selecting $\alpha_{1}=\alpha_{2}=10^{3}$ so as to make equation loss tends towards even lower order of magnitudes. Also, we show in Table 3 the error values $\epsilon$ for $u_{x}$ and $u_{y}$. We observe that the testing errors are one order higher than the training errors. However, they are still low enough to obtain physically meaningful solutions. In Fig. 10, we show the contour plots from the PINN model and FEM for testing $\mu=0.075$ MPa. We notice a high accuracy in the prediction of $u_{x}$ and $u_{y}$ from PINN. Even if stresses are not exactly matching to the FEM solution, they are of the same order of magnitude. Furthermore, in Fig. 11, we show the error $\epsilon$ as a function of time for a particular location in space. We observe again a very close match between PINN and FEM for the displacements. Barring the training time for PINN, the prediction based on PINN for new parameters $\mu$ is extremely fast, usually seconds. For the same problem where $\mu=0.075$ MPa, the FEM simulation takes approximately $20$ minutes on a standalone Intel Core i7 CPU. Hence, the PINN model easily outperforms the FEM simulation in terms of computational cost while maintaining reasonable accuracy.
@@
Figure 11: Line plots of the quantities of interest at (x=120mm,y=8mm)(x=120\ $\text${mm},\ y=8\ $\text${mm}) for material property $\mu=0.075$ MPa. Plots in (a) to (e) show the comparison of displacements and stresses obtained by PINN with respect to the FEM solutions.
@@ 5.4.2 Results and discussion (3D)
Figure 13: Contour plots of the quantities of interest in the xyxy-plane at z=5mmz=5\ mm and $t=2.29$ s for the material property $\mu=0.075$ MPa. Plots in (a) denote FEM solutions. Plots in (b) denote PINN solutions. Plots in (c) denotes the absolute error.
Figure 14: Line plots of the quantities of interest at (x=120mm,y=8mm,z=5mm)(x=120\ $\text${mm},\ y=8\ $\text${mm},\ z=5\ $\text${mm}) for material property $\mu=0.075$ MPa. Plots in (a) to (e) show the comparison of displacements and stresses obtained by PINN with respect to the FEM solutions.
@@
$0.2$ $3.5\times 10^{-3}$ $\times 10^{-3} 1.1$$1\times 10^{-3}$